what is the solution set of the inequality 5-|x+4|
Answer by stanbon(57219)

what is the solution set of the inequality
------------------------
5-|x+4| <= -3
Rearrange:
|x+4| >= 8
----
x+4 <= -8 or x+4 >= 8
---
x <= -12 or x >= 4

---
Cheers,
Stan H.